# Redefining spectral unmixing for in-vivo brain tissue analysis from hyperspectral imaging

### v2

@@ 1 Introduction
2. We show how the proposed procedure can improve targeted endmember detection methods, such as Orthogonal Subspace Projection (OSP) and Constrained Energy Minimization (CEM). During unmixing, these methods suppress contributions to the reflection spectrum from all molecules except for a selected subset of molecules of interest but can be suboptimal in extracting plausible molecular maps.
@@ 2 Methods
For our experiments, we used a publicly available HSI dataset of in-vivo brain tissue [12]. The spectral range of the used HSI system goes from 400 to 1000 $\mathrm{nm}$. Across the spectral range, 826 bands are recorded with a uniform step size and bandwidth of 2 to 3 $\mathrm{nm}$. The dataset also includes labels for four classes of pixels: Normal tissue (NT), tumor tissue (TT), blood vessels (BV), and background (BG). Over the whole dataset, images were sparsely labeled for the NT (300339 pixels), TT (21251), BV (98783) and BG class (205467), see Fig. 3 for an example of a labeled HSI image.
@@ 3 Results and Discussion
For all methods, we analysed absolute and relative molecular heatmaps. For the relative heatmaps, before performing OSP or ICEM, a reference spectrum $I_{A}^{ref}$, is subtracted from each spatial pixel. As the reference spectrum, a randomly selected pixel labeled as normal tissue was selected from each HSI. For the absolute heatmaps, we analysed the spectra $I_{A}$ directly.
3.1 OSP
Fig. 1 shows heatmaps obtained using the OSP method, with the set of endmembers described in Sec. 2.4, for an exemplary HSI image labeled as 12-1 in the Helicoid dataset.
Figure 1: Differential heatmaps for an HSI image (12-1) obtained using OSP.
For the heatmaps obtained with the OSP, Fig. 1, the following observations can be made. The tumor area displays notably lower values, e.g. for HbO2, oxCytc, which is in line with the hypoxic nature of tumor pathophysiology. Conversely, lipids exhibit increased values within the tumor area, which is again expected as lipid buildup in tumor is a known biomarker. Furthermore, the HHb heatmap effectively delineates the vascular structures of the large vessels, whereas the HbO2 heatmap contrasts smaller capillaries. These trends are consistent across multiple other cerebral HSI images in our experiments.
3.2 ICEM
Comparing the generated heatmaps for the ICEM method with different $\kappa$, the maps with $\kappa=1$ appeared least noisy and most physiologically plausible, and thus, we used the ICEM maps with $\kappa=1$ in all our experiments.
However, even for $\kappa=1$, Fig. 2, the heatmaps for the cytochromes, hemoglobin, and scattering exhibit a high similarity. All of their heatmaps equally highlight vascular structures and the tumor region, while according to our prior biological knowledge, the molecules should highlight semantically different tissue areas.
@@ 3.2 ICEM
A different pattern is observed though when we base the spectral unmixing on the pseudo-endmembers obtained from the MC simulations. Cytochromes, the metabolic molecules, clearly outline the proliferating tumor area, while hemoglobins, the main molecules of blood, contrast the vascular tree.
3.3 Quantitative Evaluation
@@ 4 Conclusion
In this study, we propose a novel spectral unmixing methodology for hyperspectral imaging of in-vivo brain tissue addressing limitations of standard linear unmixing. By using pseudo-endmembers derived from Monte Carlo simulations, our method enables a more accurate representation of light-matter interaction and enhances molecular biomarker extraction. Our results demonstrate that this approach improves the classification of brain tissue types reinforcing its potential applications in tissue characterization. While we employ a Monte Carlo simulator in our work, the methodology is applicable to other complex simulators. We thus envision that the wider optics research community will adopt this methodology, contributing to more refined and generalizable spectral imaging analysis.

### v3

@@ 1 Introduction
2. We show how the proposed procedure can improve targeted endmember detection methods, such as Orthogonal Subspace Projection (OSP) and Constrained Energy Minimization (CEM). During unmixing, these methods allow to extract contributions to the reflection spectrum from a selected subset of molecules of interest but can be suboptimal in extracting plausible molecular maps.
@@ 2 Methods
For our experiments, we used a publicly available HSI dataset of in-vivo brain tissue [12]. The spectral range of the used HSI system goes from 400 to 1000 $\mathrm{nm}$. Across the spectral range, 826 bands are recorded with a uniform step size and bandwidth of 2 to 3 $\mathrm{nm}$. The dataset also includes labels for four classes of pixels: Normal tissue (NT), tumor tissue (TT), blood vessels (BV), and background (BG). Over the whole dataset, images were very sparsely labeled for the NT (300339 pixels), TT (21251), BV (98783) and BG class (205467), see Fig. 2 for an example of a labeled HSI image.
@@ 3 Results and Discussion
For all methods, we analysed absolute and differential molecular heatmaps. For the differential heatmaps, before performing OSP or ICEM, a spectrum belonging to a randomly selected pixel is subtracted from all other spectra of the image. For the absolute heatmaps, we analysed the spectra $I_{A}$ directly. In Fig. 1, one can see an example of ICEM maps inferred from a typical HSI image. We see that the ICEM heatmaps (with $\kappa=1$ that appeared least noisy) for the cytochromes, hemoglobin, and scattering exhibit a high similarity. All of their heatmaps equally highlight vascular structures and the tumor region, while according to our prior biological knowledge, the molecules should highlight semantically different tissue areas. A different pattern is observed though when we base the spectral unmixing on the pseudo-endmembers obtained from the MC simulations. Cytochromes, the metabolic molecules, clearly outline the proliferating tumor area, while hemoglobins, the main molecules of blood, contrast the vascular tree.
@@
[…]
Next, in contrast to linear unmixing, the non-linear one is way more challenging to make work due to greater ill-posedness, no closed-form solution, and thus higher sensitivity to noise. For the Helicoid data of moderate SNR, this leads to unstable spectral unmixing, as we observed in our prior experiments. The same is found in the literature (see Fig. 6b in [16], NEBEAE method), where unstable unmixing leads to radical degradation of performance on the segmentation task. That was the reason that brought us to develop the proposed methodology which is robust against the typical obstacles for non-linear unmixing.
@@ 4 Conclusion
In this study, we propose a novel spectral unmixing methodology for hyperspectral imaging of in-vivo brain tissue addressing limitations of standard unmixing. By using pseudo-endmembers derived from Monte Carlo simulations, our method enables a more accurate representation of light-matter interaction and enhances molecular biomarker extraction. Our results demonstrate that this approach improves the classification of brain tissue types reinforcing its potential applications in tissue characterization. While we employ a Monte Carlo technique in our work, the methodology is applicable to other physics simulators. We thus envision that the wider optics research community will adopt this methodology, approbating various complex light-matter modeling techniques.
Finally, we do not claim clinical readiness for tumor classification with the current F1 score accuracy by SOTA methods (the classification analysis was just one of several ways to quantitatively assess the utility of the extracted biomarkers by the proposed method). But we do believe that to achieve it, improvements in the methodology and improvements in the quality of ground truth annotation have to go hand in hand to mutually benefit from each other. Better model predictions will help to better annotate the images by experts in a semi-automated fashion, and vice versa, more and better annotated data will lead to superior models.